# Twist, splay, and uniform domains in ferroelectric nematic liquid crystals

### v1

@@ IV.2 Pie slice domains
where we have assumed that $\partial_{\theta}\phi\ll 1$ is negligibly small throughout most of the sample. Regions of opposite (cancelling) charge are created by alternating between $\phi(\theta)=0,\pi$, corresponding to $\mathbf{P}$ parallel or anti-parallel, respectively, to $\hat{\mathbf{r}}$.
We consider a single-mode approximation so that $\cos[\phi(\theta)]\approx\cos(n_{\theta}\theta)$ and $2n_{\theta}$ is the total number of polarization domains [pie slices in Fig. 7(a)]. Then, in cylindrical coordinates, the screened Coulomb potential is given by
$\frac{e^{-\kappa|\mathbf{r}-\mathbf{r}^{\prime}|}}{|\mathbf{r}-\mathbf{r}^{\prime}|}$ $=\frac{2}{\pi}\sum_{n=0}^{\infty}(2-\delta_{n})\cos[n(\theta-\theta^{\prime})]$
×$\times\int_{0}^{\infty}\mathrm{d}k\,I_{n}(x_{<})K_{n}(x_{>})\cos[k(z-z^{\prime})],$ (25)
where we have the Kronecker delta $\delta_{n}$ ($\delta_{n}=1$ if $n=0$ and $\delta_{n}=0$, otherwise). Also, $x_{<,>}=\sqrt{k^{2}+\kappa^{2}}\,r_{<,>}$ and $r_{<}$ ($r_{>}$) is the smaller (larger) of the polar distances $r$ and $r^{\prime}$. Substituting $\rho(\theta)$ from Eq. (24) and Eq. (25) into Eq. (1) yields (after some algebra and an identity for the integral of a single modified Bessel function of the first kind $I_{\nu}(z)$ Gradshteyn and Ryzhik (2007))
$F_{\rho}$ $=\frac{32\pi^{2}h^{3}P_{0}^{2}}{\epsilon\epsilon_{0}}\sum_{m=0}^{\infty}(-1)^{m}\int_{0}^{\infty}\mathrm{d}u\,\frac{\sin^{2}(\frac{u}{2})}{u^{2}(u^{2}+h^{2}\kappa^{2})}$
×$\times\,\int_{0}^{\frac{R}{h}\sqrt{u^{2}+h^{2}\kappa^{2}}}\mathrm{d}v\,\,I_{2m+1+n_{\theta}}(v)K_{n_{\theta}}(v),$ (26)
@@
So, what about the elastic and anchoring energy? The total length of domain wall is $2n_{\theta}R$ so that the total energy of the domain walls is $F_{\mathrm{dw}}=2n_{\theta}Rf_{\pi}$, with $f_{\pi}$ the linear energy density of the domain walls (e.g., either the solitonic wall or the disclination line pair). Combining the elastic and electrostatic contributions, we find that the total free energy $F=F_{\rho}+F_{\mathrm{dw}}$ of the “pie-slice” configuration with $n_{\theta}$ cuts (i.e., $2n_{\theta}$ pie slices) reads
@@
where $F_{0}=2Rf_{\pi}$ is a characteristic energy, and $\eta_{\theta}=2\pi h/\lambda_{\mathrm{dw}}$ is the optimal number of sectors without screening $\eta_{\theta}=n_{\theta}^{*}(\kappa\rightarrow 0)$, and $\Xi_{n_{\theta}}(z)\equiv\int_{0}^{z}I_{n_{\theta}+1}(z)K_{n_{\theta}}(z)\,\mathrm{d}z$. The integrations can be evaluated numerically. The plot of the total free energy in Eq. (27) is shown in Fig. 7(b). We see that the total free energy exhibits a minimum at a non-zero value of $n_{\theta}$, analogously to the plot in Fig. 6(b) for the stripe domains. We thus find that the pie-slice configuration, such as the one shown schematically in Figs. 7(a) and in a micrograph in Fig. 2(a,b), is energetically favorable.
At this point, we can look for the behavior of the inside the integrals in Eq. (26) for large $n_{\theta}$ to get a better sense of the Coulomb energy contribution. This is possible via asymptotic expansions Sidi and Hoggan (2011). There are two cases to consider: $\kappa R\ll n_{\theta}$ (weak screening) and $\kappa R\gg n_{\theta}$ (strong screening). We find that, for thin cells compared to the domain size, $h\ll R$,
@@
We see that in both cases, this Coulomb energy decreases with increasing $n_{\theta}$, in contrast to the elastic and anchoring energies which will increase proportionally to $n_{\theta}$ due to the domain wall formation. One may compare the energetic contribution on the first line of Eq. (28) to Eq. (8), where the dominant contribution to the electrostatic energy scales as $F_{\rho}\propto I_{1}(\alpha_{0})K_{1}(\alpha_{0})\propto 1/k_{z}$ for sufficiently small $\kappa$. We see here that we get a similar behavior, in that $F_{\rho}\propto 1/n_{\theta}$ in Eq. (28). In other words, in both cases, the electrostatic interaction generates an energetic contribution inversely proportional to the “wave number” of the spatial modulation.
We can minimize the total free energy $F$ to find a pretty simple approximate result:
@@
where $\lambda_{\mathrm{dw}}=\sqrt{2f_{\pi}\epsilon\epsilon_{0}}/P_{0}$ is the previously-mentioned characteristic length and the large versus small screening conditions have to be checked self-consistently. The value $2n_{\theta}^{*}$ should give the “optimal” number of sectors (i.e., $\pm$ polarization domains).
This description, however, is incomplete as we also need to consider what happens very far from the center of the radial aster where the anchoring looks uniform along the local radial direction. Here, we expect to again find parallel striped domains, just as in the stripe domain case considered in the previous subsection: The sample should break up into stripes along the radial direction with characteristic wavelength $\lambda_{x}^{*}$. We can use the analysis of the striped pattern and imagine that the circumference $2\pi R$ breaks up into segments with wavelength $\lambda_{x}^{*}$. This corresponds to an optimal number (in the low screening limit $\kappa\lambda_{\mathrm{dw}}\ll\sqrt{h/R}$) of
$n_{\theta}^{\mathrm{stripe}}\approx\frac{2\pi R}{\lambda_{x}^{*}}\approx 1.2\,\dfrac{R}{\,\lambda_{\mathrm{dw}}}\sqrt{\frac{h}{L}},$ (30)
where $L$ is a large dimension (i.e., the full sample length). It is also worth noting that we have not considered the energetic cost of nucleating the domain walls near the aster center where the domain walls get close to each other. This would generate additional elastic distortion, so we might expect that Eq. (29) generally overestimates the number of sectors.
V Quantitative Comparisons
@@
Figure 8: (a) TD stripe wavelength $\lambda_{x}^{*}\approx 2\langle w_{\mathrm{stripe}}\rangle$ for various cell thicknesses $h$ calculated form the average stripe width across a sample with unidirectional, bipolar anchoring. The error bars show the standard deviation over observed stripes. The red solid curve is the wavelength found by minimizing Eq. (20) over $\kappa\lambda_{x}$ for a fixed $\kappa\lambda_{\mathrm{dw}}=0.01$ and $\lambda_{D}=\kappa^{-1}=10~{}\mu\mathrm{m}$. The blue dashed line is the approximation in Eq. (21). (b) The number of “pie-slice” sectors (SDs) $2n_{\theta}^{*}$ counted out at a distance of $R=103.6~{}\mu\mathrm{m}$ from the defect center of a cell pre-patterned with a $+1$ “aster” defect. The red line corresponds to the sector number evaluated by minimizing the free energy in Eq. (27) with $\lambda_{D}=\kappa^{-1}=0.18~{}\mu\mathrm{m}$ and $\lambda_{\mathrm{dw}}=0.1~{}\mu\mathrm{m}$. The blue dashed line is the approximation in Eq. (29) (second line). The orange dotted line is the “stripe” approximation in Eq. (30) with $\lambda_{\mathrm{dw}}=0.1~{}\mu\mathrm{m}$, $\lambda_{D}=10~{}\mu\mathrm{m}$ and $L_{y}=1~{}\mathrm{cm}$
To compare these theoretical results to experiment, it is necessary to make some estimates for the domain wall energy density $f_{\pi}$ which determines the characteristic length $\lambda_{\mathrm{dw}}$. A key question is how $f_{\pi}$ depends on the cell thickness $h$. For simplicity, we will assume that the walls are the disclination pairs, so that $f_{\pi}\approx 2K\sim 10-1000~{}\mathrm{pN}$ is independent of the thickness $h$. Then, estimating that $P_{0}\approx 0.044~{}\mathrm{C}/\mathrm{m}^{2}$ for DIO Nishikawa et al. (2017b) and $\epsilon\approx 10-100$, we find the range $\lambda_{\mathrm{dw}}=(2f_{\pi}\epsilon\epsilon_{0})^{1/2}P_{0}^{-1}\sim 3~{}\mathrm{nm}-30~{}\mathrm{nm}$. This is a small length corresponding to the structure of the domain wall.
We also need to estimate the screening length $\lambda_{D}=\kappa^{-1}$. This is not necessarily constant throughout the sample as different amounts of free charge may accumulate throughout the sample, so the bound charges may not be uniformly cancelled. For example, in the stripe case (UDs and TDs), the screening effects would be most important near the domain or cell boundary. On the other hand, for the radial slices (SDs), the screening near the $+1$ defect would be more relevant. Moreover, as discussed previously, the screening depends on the sample preparation protocol, such as the cooling rate. Thus, we might expect different values of $\kappa$ depending on the sample considered.
For the stripe patterns (TDs), we measure the stripe widths $w_{\mathrm{stripe}}$ across a cell over a distance of about $500~{}\mu\mathrm{m}$ [see Fig. 1(a-d)]. The average $\langle w_{\mathrm{stripe}}\rangle$ and standard deviation $\sqrt{\langle w_{\mathrm{stripe}}^{2}\rangle-\langle w_{\mathrm{stripe}}\rangle^{2}}$ for various cell thicknesses between $2~{}\mu\mathrm{m}$ and $10~{}\mu\mathrm{m}$ are shown as the data points and error bars, respectively, in Fig. 8(a). To compare to theory, we minimize Eq. (20) with respect to $\kappa\lambda_{x}$ while varying the parameters $\kappa$ and $\lambda_{\mathrm{dw}}$. The red dashed line in Fig. 8(a) corresponds to $\kappa\lambda_{\mathrm{dw}}=0.01$ and $\lambda_{D}=\kappa^{-1}=10~{}\mu\mathrm{m}$, with $L_{y}=1~{}\mathrm{cm}$ the full sample size. This is a relatively large screening length and a correspondingly large $\lambda_{\mathrm{dw}}\approx 100~{}\mathrm{nm}$. This suggests that the domain walls between polarization domains cost more energy than our simple estimate suggests and that the depolarization fields at the sample boundaries are not strongly screened. The zero screening prediction $(\kappa\rightarrow 0)$ is also shown in Fig. 8(a).
For the pie-slice domains in the $+1$ aster defect cell, we count the number of domains out at a radius $R\approx 100~{}\mu\mathrm{m}$ away from the center of the defect. We then compare to the theoretical result for $2n_{\theta}^{*}$ given by Eq. (29). We find that the $\kappa R\ll n_{\theta}^{*}$ result cannot describe the data without introducing anomalously large values of either the elastic constant $K$ or the dielectric constant $\epsilon$. Instead, the data is more consistent with large screening $\kappa R\gg n_{\theta}^{*}$. Using $\lambda_{\mathrm{dw}}=100~{}\mathrm{nm}$ as for the stripe case, we find that $\lambda_{D}=\kappa^{-1}=0.13~{}\mu\mathrm{m}$ does a reasonable job describing the data, as shown by the red curve [calculated by numerically minimizing the free energy in Eq. (27)] and the blue curve [calculated using the second line in Eq. 29, shown in Fig. 8(b)]. We also checked whether the approximation in Eq. (30) yields reasonable results using the same value $\lambda_{D}=10~{}\mu\mathrm{m}$ (and $L_{y}=1~{}\mathrm{cm}$) used for the stripes. We see that this approximation overestimates the number of sectors, suggesting that the radial sector pattern formation indeed comes from unscreened bound charge generated by the $+1$ defect.
In summary, the quantitative comparisons for the pie slice (SD) and stripe domain (TD) patterns yield rather different results for the Debye screening length $\lambda_{D}$. This could be due to the different nature of the patterns: the stripes are generated due to uncompensated bound charges at the sample boundaries while the radial pie slices form due to a non-vanishing bound charge density at the $+1$ defect center. It is likely that the free ions are able to screen the charge better at the defect center. It would be interesting to systematically vary the free ion concentration in the samples in order to get a better comparison to the theoretical results, which depend sensitively on $\lambda_{D}$.
VI Conclusions
We have demonstrated that the $\mathrm{N}_{\mathrm{F}}$ spontaneously forms a spatial modulation due to the competition between Coulomb and elastic interactions. The long-range Coulomb interactions generate a large energetic penalty for regions with uniform $\mathbf{P}$. This can come from either bound charge in the bulk of the system due to a nonvanishing divergence $\nabla\cdot\mathbf{P}$ or due to uncompensated charges at the edges of domains. Overall, the Coulomb interactions tend to create regions of opposing $\mathbf{P}$ directions.

### v2

@@ Results
where we have assumed that $\partial_{\theta}\phi\ll 1$ is negligibly small throughout most of the sample. Regions of opposite (cancelling) charge are created by alternating between $\phi(\theta)=0,\pi$. Note that this approximation will break down close to the $+1$ defect (the small $R$ region) where all of the domain tips collide. The defect structure there is complex and, even in the absence of domains, the $+1$ defect splits into two $+1/2$ defects joined by a domain wall, as discussed previously (see Fig. 2h). We thus consider intermediate values of $R\approx 100~{}\mu\mathrm{m}$ when comparing to experiments in the next section.
We perform a single-mode approximation so that $\cos[\phi(\theta)]\approx\cos(n_{\theta}\theta)$. Then, in cylindrical coordinates, the screened Coulomb potential is given by
$\frac{e^{-\kappa|\mathbf{r}-\mathbf{r}^{\prime}|}}{|\mathbf{r}-\mathbf{r}^{\prime}|}=\frac{2}{\pi}\sum_{n=0}^{\infty}(2-\delta_{n})\cos[n(\theta-\theta^{\prime})]\ \int_{0}^{\infty}\mathrm{d}k\,I_{n}(x_{<})K_{n}(x_{>})\cos[k(z-z^{\prime})],$ (24)
where $\delta_{n}=1$ if $n=0$ and $\delta_{n}=0$, otherwise. Also, $x_{<,>}=\sqrt{k^{2}+\kappa^{2}}\,r_{<,>}$ and $r_{<}$ ($r_{>}$) is the smaller (larger) of the polar distances $r$ and $r^{\prime}$. Substituting $\rho(\theta)$ from Eq. (23) and Eq. (24) into Eq. (1) yields (after some algebra and an identity for the integral of a single modified Bessel function of the first kind $I_{\nu}(z)$ [38])
$F_{\rho}=\frac{32\pi^{2}h^{3}P_{0}^{2}}{\epsilon\epsilon_{0}}\sum_{m=0}^{\infty}(-1)^{m}\int_{0}^{\infty}\mathrm{d}u\,\frac{\sin^{2}(\frac{u}{2})}{u^{2}(u^{2}+h^{2}\kappa^{2})}\,\int_{0}^{\frac{R}{h}\sqrt{u^{2}+h^{2}\kappa^{2}}}\mathrm{d}v\,\,I_{2m+1+n_{\theta}}(v)K_{n_{\theta}}(v),$ (25)
@@
The total length of domain wall in the system is $2n_{\theta}R$ so that the total elastic cost is $F_{\mathrm{dw}}=2n_{\theta}Rf_{\pi}$, with $f_{\pi}$ the domain wall linear energy density (e.g., either the soliton, the domain wall, or the disclination line pair). We find that the total free energy $F=F_{\rho}+F_{\mathrm{dw}}$ of the pie-slice configuration with $n_{\theta}$ cuts (i.e., $2n_{\theta}$ pie slices) reads
@@
where $F_{0}=2Rf_{\pi}$ is a characteristic energy, $\eta_{\theta}=2\pi h/\lambda_{\mathrm{dw}}$ is the optimal number of sectors without screening $\eta_{\theta}=n_{\theta}^{*}(\kappa\rightarrow 0)$, and $\Xi_{n_{\theta}}(z)\equiv\int_{0}^{z}I_{n_{\theta}+1}(z)K_{n_{\theta}}(z)\,\mathrm{d}z$. The parameter $\lambda_{\mathrm{dw}}\equiv\sqrt{2f_{\pi}\epsilon\epsilon_{0}}/P_{0}$ depends on $f_{\pi}$ just as in the stripe case. The plot of Eq. (26) (with numerical evaluation of the integrals) is shown in Fig. 7b. The total free energy exhibits a minimum at a non-zero value of $n_{\theta}$.
We consider two cases: $\kappa R\ll n_{\theta}$ (weak screening) and $\kappa R\gg n_{\theta}$ (strong screening). We find [using asymptotic expansions of the integrals in Eq. (25) [39]] that, for thin cells compared to the domain size, $h\ll R$,
@@
In both cases, this Coulomb energy decreases with increasing $n_{\theta}$, in contrast to the elastic and anchoring energies which will increase proportionally to $n_{\theta}$. We minimize the total free energy $F$ to find the approximate result:
@@
where the large versus small screening conditions have to be checked self-consistently. The value $2n_{\theta}^{*}$ gives the “optimal” number of sectors (i.e., ±$\pm\hat{\mathbf{r}}$ polarization domains).
Quantitative comparisons between models and experiments
@@
Figure 8: Quantitative comparison. Experimental (points) and theoretical predictions (dashed lines) for the number of domains as a function of cell thickness $h$. In the theoretical models, we fix $\lambda_{D}=\kappa^{-1}=10~{}\mu\mathrm{m}$ throughout. a TD stripe wavelength $\lambda_{x}^{*}\approx 2\langle w_{\mathrm{stripe}}\rangle$ calculated form the average stripe width across a sample with unidirectional, bipolar anchoring. The error bars show the standard deviation over observed stripes. The dashed curves are the wavelengths found by minimizing Eq. (19) over $\kappa\lambda_{x}$ for various models of the domain wall cost, which impacts the characteristic length $\lambda_{\mathrm{dw}}=\sqrt{2\epsilon\epsilon_{0}f_{\pi}}/P_{0}$. We set $\lambda_{\mathrm{dw}}=0.12~{}\mu\mathrm{m}$ constant for the disclination line case (red curve). We set $\lambda_{\mathrm{dw}}=(0.0608~{}\mu\mathrm{m}^{3/4})h^{1/4}$ for the soliton (green), and $\lambda_{\mathrm{dw}}=(0.0437~{}\mu\mathrm{m}^{1/2})h^{1/2}$ for the fixed-width wall (purple). b The number of pie-slice sectors (SDs) $2n_{\theta}^{*}$ counted out at a distance of $R=104~{}\mu\mathrm{m}$ from the defect center of a cell pre-patterned with a $+1$ “aster” defect. The theory curves are found by minimizing the free energy in Eq. (26) with respect to $n_{\theta}$. We again consider the disclination model with $\lambda_{\mathrm{dw}}=1.5~{}\mu\mathrm{m}$ constant, a soliton with $\lambda_{\mathrm{dw}}=(1.4~{}\mu\mathrm{m}^{3/4})h^{1/4}$, and the fixed-width wall with $\lambda_{\mathrm{dw}}=(1.3~{}\mu\mathrm{m}^{1/2})h^{1/2}$. The error bars are estimated from the uncertainty in the domain count in the sample.
A key question is how the domain wall energy density $f_{\pi}$ depends on the cell thickness $h$. One possibility is that the walls are disclination pairs, so that $f_{\pi}\sim 2K\sim 10-1000~{}\mathrm{pN}$ is independent of the thickness $h$ and is proportional to some combination of elastic constants, here indicated as $K$. However, for smaller thicknesses $h$, it is likely that the elastic distortion at the wall extends through the entire cell thickness $h$ and contributes meaningfully to the energy of the domain wall. In this case, we can consider other scalings such as $f_{\pi}\propto\sqrt{h},h$ for the soliton and fixed-width domain wall, respectively. These various scalings, then, will change how the characteristic length $\lambda_{\mathrm{dw}}=(2f_{\pi}\epsilon\epsilon_{0})^{1/2}P_{0}^{-1}$ varies with $h$, with $\lambda_{\mathrm{dw}}\propto h^{0},h^{1/4},h^{1/2}$ for the disclinations, soliton, and fixed-width domain wall, respectively. These three possibilities are shown in Fig. 8 for both the stripe patterns and the pie-slice sector results, which we consider in more detail below.
The screening length $\lambda_{D}=\kappa^{-1}$ is not necessarily constant throughout the sample. For example, in the stripe case (UDs and TDs), the screening effects would be most important near the domain or cell boundary. On the other hand, for the radial slices (SDs), the screening near the $+1$ defect would be more relevant. The value of $\kappa$ may also change with the cell thickness $h$, as ions may penetrate the sample less when the thickness $h$ is less than the screening length $\kappa^{-1}$. We will fix $\kappa=0.1~{}\mu\mathrm{m}^{-1}$ in our analysis, which seems to work well for the samples considered here. This is a relatively large screening length ($\lambda_{D}=\kappa^{-1}=10~{}\mu\mathrm{m}$), but not necessarily large enough that we can take the zero screening limit.
For the stripe patterns (TDs), we measure the stripe widths $w_{\mathrm{stripe}}$ across a cell over a distance of about $500~{}\mu\mathrm{m}$ (see Fig. 1a-d). The average $\langle w_{\mathrm{stripe}}\rangle$ and standard deviation $\sqrt{\langle w_{\mathrm{stripe}}^{2}\rangle-\langle w_{\mathrm{stripe}}\rangle^{2}}$ for various cell thicknesses are shown as the data points and error bars, respectively, in Fig. 8a. To compare to theory, we minimize Eq. (19) with respect to $\lambda_{x}$ to find the optimal value $\lambda_{x}^{*}$ for various values of $h$ and fixed $\lambda_{D}=\kappa^{-1}=10~{}\mu\mathrm{m}$ and full sample size $L=1~{}\mathrm{cm}$. We also vary the value of $\lambda_{\mathrm{dw}}$ to find a favorable match to the experimental data, trying different possibilities for the $h$-dependence: The red dashed line in Fig. 8a corresponds to a constant $\lambda_{\mathrm{dw}}=0.12~{}\mu\mathrm{m}$, which would be consistent with a domain wall consisting of disclination lines. The green dashed line in Fig. 8a corresponds to a solitonic domain wall with $f_{\pi}=f_{\pi}^{\mathrm{soliton}}\propto\sqrt{h}$, for which we set $\lambda_{\mathrm{dw}}=(0.0608~{}\mu\mathrm{m}^{3/4})h^{1/4}$, a scaling which compares favorably with the data for intermediate thicknesses $h$ between 2 and 8 microns. Finally, if the domain wall elastic distortion extends across the entire cell thickness and we have $f_{\pi}=f_{\pi}^{\mathrm{wall}}\propto h$, then we find that we get good agreement with the data for $\lambda_{\mathrm{dw}}=(0.0437~{}\mu\mathrm{m}^{1/2})h^{1/2}$, as shown by the purple dashed line in Fig. 8a. Note that in this case, the stripe wavelength $\lambda_{x}^{*}$ does not change with the thickness $h$, which is consistent with the experimental results. We may thus conclude that, in our samples, the domain walls between the polarization domains likely have an energetic cost that scales proportionally to the cell thickness $h$.
For the pie-slice domains in the $+1$ aster defect cell, we count the number of domains out at a radius $R=104~{}\mu\mathrm{m}$ away from the center of the defect. We then compare to the theoretical result for $2n_{\theta}^{*}$ given by the minimum of the energy in Eq. (26) with respect to $n_{\theta}$. The red dashed curve in Fig. 8b has a constant $\lambda_{\mathrm{dw}}=1.5~{}\mu\mathrm{m}$, corresponding to disclinations as the domain walls. This clearly underestimates the energetic cost of the domain walls, especially for $h>4~{}\mu\mathrm{m}$. On the other hand, if we consider the solitonic wall (green curve in Fig. 8b) and set $\lambda_{\mathrm{dw}}=(1.4~{}\mu\mathrm{m}^{3/4})h^{1/4}$, we get a better match to the data over a larger range of $h$. Finally, for the fixed-width domain walls with energies proportional to $h$, we find a good match to the data with $\lambda_{\mathrm{dw}}=(1.3~{}\mu\mathrm{m}^{1/2})h^{1/2}$, shown by the purple dashed curve in Fig. 8b. Thus, like in the striped pattern case, we conclude here that the cost of the domain walls between polarization domains likely scales linearly with the cell thickness $h$. However, we note that we have fixed $\kappa=0.1~{}\mu\mathrm{m}^{-1}$ to a constant and $\kappa$ may well be $h$-dependent. Moreover, the $h$-dependence of $\lambda_{\mathrm{dw}}$ might be more complex than the cases we have considered here. Nevertheless, we note the good agreement between the theoretical results with fixed-width domain wall energy $f_{\pi}=f_{\pi}^{\mathrm{wall}}\propto h$ and the experimental data for both the stripe wavelength calculation and the calculated number of pie slices in Fig. 8a,b, respectively.
In summary, the theoretical results for the stripe domain (TD) and pie slice (SD) patterns match favorably with experimental data, as shown in Fig. 8a,b, respectively. The two cases seem to have different values for the characteristic domain wall length $\lambda_{\mathrm{dw}}$, with the SD domains consistent with a larger (approximately 10 to 30-fold) value of $\lambda_{\mathrm{dw}}$. This could be due to different values of the effective dielectric constant $\epsilon$, or it could be that the two cases have different values of $\kappa$, which we kept constant here at $\kappa=0.1~{}\mu\mathrm{m}^{-1}$ for both cases. Changing the value of $\kappa$ would result in different values of $\lambda_{\mathrm{dw}}$ when matching to experimental data. Variations in $\kappa$ and $\lambda_{\mathrm{dw}}$ could be due to the different nature of the patterns: the stripes are generated due to uncompensated bound charges at the sample boundaries while the radial pie slices form due to a non-vanishing bound charge density at the $+1$ defect center. Free ions may be able to screen the charge better at the defect center, leading to larger $\kappa$ values and, similarly, larger values of $\lambda_{\mathrm{dw}}$ (which is proportional to $\sqrt{\epsilon}$) for the SD case relative to the TD pattern.
Discussion
The long-range Coulomb interactions for regions with uniform $\mathbf{P}$ generate a large electrostatic energy penalty at region boundaries. Overall, the Coulomb interactions tend to create regions of opposing $\mathbf{P}$ directions. On the other hand, reorientations of the polarization direction incur elastic energy costs. We have demonstrated that the competition between these two effects creates a variety of domain patterns that depend strongly on the cell thickness and the ionic content.